solve the (square root of x+4)+8=0
Found 2 solutions by haileytucki, ovie27:
Answer by haileytucki(390)

~((x+4))+8=0
Since 8 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 8 from both sides.
~((x+4))=-8
To remove the radical on the left-hand side of the equation, square both sides of the equation.
(~((x+4)))^(2)=(-8)^(2)
Simplify the left-hand side of the equation.
(x+4)=(-8)^(2)
Squaring an expression is the same as multiplying the expression by itself 2 times.
(x+4)=((-8)(-8))
Multiply -8 by -8 to get 64.
(x+4)=((64))
Remove the parentheses around the expression 64.
(x+4)=(64)
Remove the parentheses around the expression 64.
(x+4)=64
Remove the parentheses around the expression x+4.
x+4=64
Since 4 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 4 from both sides.
x=-4+64
Add 64 to -4 to get 60.
x=60
Verify each of the first set of solutions by substituting them into the original equation and solving. In this case, none of the solutions are valid.
No Solution

Answer by ovie27(16)


sqrt(x+4) + 8 = o
bring the 8 to the other side:
sqrt(x+4) = -8
square both sides to get rid of the square root:
(x+4) = 64
bring the 4 to the other side:
x = 60
The thing is when you plug in the 60 into the equation, and you take the square root of 64, it can be + or - 8. if it is negative, then the equation works, and if it is positive, then the answer is not 0, it is 16:
sqrt(64) + 8 =0
8 + 8 = 16 which does not equal 0, so
we have a restriction to this problem: sqrt(64) cannot be positive.
we get -8 + 8 = 0.